A fire hose ejects a stream of water at an angle of 30.3° above the horizontal. The water leaves the nozzle with a speed of 20.0 m/s. Assuming that the water behaves like a projectile, how far from a building should the fire hose be located to hit the highest possible fire?
1
Expert's answer
2011-11-17T08:30:43-0500
a= 30 - angle
v=20 m/s
t- time when one projectile of water will rich middle of trajectory
v*sina= gt
t=20*0.5/10= 1sec

S=v*cosa*t=sqrt(3)/2*20=10sqrt(3)
Answer: 10sqrt(3)